# Entering the overcritical regime of nonlinear Breit-Wheeler pair production in collisions of bremsstrahlung $\gamma$-rays and superintense, tightly focused laser pulses

### v1

@@ III.1 Comparison between S-Model and D-Model at moderately high field strengths
Figure 3: Comparison between the D-Model (solid) and S-Model (dashed) of differential number of created pairs from different bremsstrahlungs photon energies at $\xi=70$ and $E_{0}=2.5$ GeV.
Figure 3 shows the differential number of created pairs (per radiating electron) in dependence on the normalized bremsstrahlung-photon energy. The set of parameters of a planned experiment on strong-field Breit-Wheeler pair production [13, 33] is considered here: $E_{0}=2.5$ GeV, $\xi=70$ (corresponding to an intensity of $I\approx 10^{22}$ W/cm2), and $\tau=30$ fs. The highest value of the quantum nonlinearity parameter resulting in this scenario is $\kappa\approx 2$ for $f\approx 1$. As the figure displays, the contribution to the pair production yields grows monotoneously with increasing $f$. Up to $f\approx 0.5$ (corresponding to $\kappa\approx 1$), the predictions of the S- and D-models closely agree with each other, but start to slightly deviate afterwards. At the bremsstrahlung endpoint $f\approx 1$, the S-model overestimates the pair yields by about 20%.
@@
Apart from the value of the quantum nonlinearity parameter $\kappa$, we emphasize that also the duration of the interaction influences the relation between S- and D-model predictions. Let us demonstrate this by a simple estimate. In the tunneling-like regime of strong-field Breit-Wheeler pair production at $\kappa\ll 1$, the process rate is given by the analytical expression $R\sim\frac{\alpha m^{2}\kappa}{\omega^{\prime}}\,{\rm e}^{-8/(3\kappa)}$. This compact formula still holds approximately for $\kappa\sim 1$ (see, e.g., Fig.5 in [33]). Taking $\kappa=2$ and $\omega^{\prime}=2.5$ GeV, the rate becomes $R\approx 10^{-7}m$. As a consequence, the time-integrated rate remains smaller than unity for laser pulse durations $\tau\lesssim 10$ fs. Therefore, the S-model is applicable for few-cycle optical laser pulses with duration below 10 fs in this case.
@@ III.2 Pair production in the overcritical field regime
In this subsection, we utilize our D-Model and apply Eq. (15) to obtain several important insights into the Breit-Wheeler pair creation process in the overcritical regime with $1<\kappa\lesssim 30$. In contrast to the previous Sec. III.1, here the attention is put on high laser intensities above $10^{22}\ \mathrm{W/cm^{2}}$. These values are achievable in the foreseeable future at the laser facilities such as ELI-NP in Romania, where the investigation of the Breit-Wheeler pair production is firmly embedded into the planning [15]. In the following, we shall therefore use a laser pulse duration of $\tau=20$ fs, as a typical value envisaged at this facility.
@@ III.2.1 Spectrally resolved pair yields
Figure 5: Differential number of created pairs from different bremsstrahlung photon energies at incident electron energy $E_{0}=2.5$ GeV for different values of the laser intensity parameter $\xi\in\{70,100,150,200,250,300\}$ as indicated using the D-model.
@@
By utilizing the D-model we can analyse the contribution to the number of created pairs from different $\gamma$-photon energies within the bremsstrahlung spectrum. These contributions are depicted in Fig. 5 for different values of $\xi$ between 70 and 300 at a fixed incident electron energy of $E_{0}=2.5$ GeV. As we go to high values of $\xi$ we see a clear contrast in the distributions compared to the case in Fig. 3. For $\xi=70$, we observe as before a strictly monotonic increase in the number of created pairs as $f$ increases. In contrast, for higher values of $\xi$, a pronounced maximum emerges, indicating a peak contribution at specific $\gamma$-photon energies. For $\xi=150$, the curve shows only a slight bend, while for $\xi=200$, we observe a global maximum at $f\approx 0.48$. Furthermore there is a maximum at $f\approx 0.36$ for $\xi=250$ and a very high peak at $f\approx 0.28$ for $\xi=300$. The formation of those maxima results from three effects which influence the creation of pairs. Firstly, for the parameters under consideration, the local rate of pair creation still approaches to an exponential dependence and thus increases steeply with $\kappa$, which is proportional to the product of $\xi$ and $f$. Thus, for $\xi=70$, the exponential rate damping is still strong and therefore the number of pairs grows as $f$ increases, whereas for high values of $\xi$, the field dependence of the rate becomes much flatter (see also Fig. 8 below). Secondly, the number of produced bremsstrahlung photons is much higher for low values of $f$, as depicted in Fig. 2, which starts playing a crucial role in the pair production when the rate growth has left the very steep exponential region and entered into the flatter portion. Lastly, the exponential decay of the bremsstrahlung photons in the D-model is much more pronounced for high values of $\kappa$, which also favors low values of $f$ for the creation of pairs. It is therefore, that we observe the emergence of a maximum at lower values of $f$, because the increase of the rate with higher values of $f$ is counteracted and eventually overcompensated by the falling distribution of bremsstrahlung photons and their decay for high values of $\xi$.
Figure 6: Contributions to the number of created pairs from different bremsstrahlung photon energies at $\xi=250$ obtained from the D-model. The red, blue and black curves (in this order) represent different values of the incident electron energy $E_{0}\in\{2.5,5,10\}$ GeV. The dashed line represents the result for $E_{0}=2.5$ GeV predicted by the S-model.
@@
A similar observation can be made when we compare the impact of the energy of bremsstrahlungs photons for different values of the incident electron energy. In Fig. 6 the contribution to the number of created pairs is plotted for $E_{0}=2.5$ GeV (red), $E_{0}=5$ GeV (blue) and $E_{0}=10$ GeV (black) at fixed $\xi=250$. A global maximum in the number of created pairs arises at $f\approx 0.35$ for $E_{0}=2.5$ GeV, $f\approx 0.185$ for $E_{0}=5$ GeV, and the most pronounced maximum for $E_{0}=10$ GeV at $f\approx 0.0975$.
@@
For comparison, the results predicted by the S-Model for $E_{0}=2.5$ GeV are represented by the dashed red line. It agrees with the D-model result up to $f\approx 0.15$, corresponding to $\kappa\approx 1.1$, but sizably overestimates the yield for higher bremsstrahlung photon energies. In particular, no maximum emerges along the S-Model curve. The S-model results for the higher electron energies are not shown in Fig. 6 because they would deviate even more strongly from the proper pair yields obtained from the D-model. We note that these curves would feature maxima at low $f$ values because of the counteracting trends between the increase of the rate and the decrease of the bremsstrahlung spectrum. An additional effect contained in the D-model description is the attenuation of the $\gamma$-beam. In result, the maxima of the blue and black line are located at smaller $f$ values than one would expect based on an S-model consideration.
@@
As an illustration, Fig. 7 depicts the values of the spacetime dependent quantum nonlinearity parameter $\kappa(t,\rho,z,f)$ at $\rho=0$ for 10 GeV incident electrons and the integration region given in Eq. (8). In the upper panel the local values of $\kappa$ for very small photon energy with $f=0.05$ is exposed. Here, the number of created particles is expected to be low as $\kappa$ is of the order of one. On the other hand, for $f=0.1$ the values presented in the middle panel overpass the critical value of $\kappa\approx 1$ and the expected pair yield per $\gamma$-photon is significantly increased, as the heavy exponential damping characteristic for the tunneling-like regime at $\kappa\ll 1$ is largely softened here, so that the course of the pair production rate is distinctly flattened. This observation combined with the high number of produced bremsstrahlung photons (see the graph in Fig. 2) is responsible for the emergence of the maximum in Fig. 6. Lastly, the quantum nonlinearity parameter for $f=0.2$ is shown in the lower panel of Fig. 7. Here, $\kappa\approx 6$ is reached leading to a high rate of the particle creation. However, as the number of bremsstrahlung $\gamma$’s is much lower than for $f=0.1$ (see Fig. 2) and their attenuation in the process is stronger, the resulting outcome for particle production is reduced.
@@ III.2.2 Total pair yields
Figure 8: Dependence of the number of created pairs on the laser intensity for various incident electron energies, color-coded with red, blue and black solid lines for $E_{0}=2.5$, 5, and 10 GeV using the D-model. The dashed line represents the result for $E_{0}=2.5$ GeV given by the S-model.
@@
In Fig. 8 the total number of pairs produced by a single radiating electron is considered in the intensity range $\xi\in[50,300]$. Moreover, three different values of the incident electron energy $E_{0}\in\{2.5,5,10\}$ GeV are studied and depicted as red, blue and black solid lines, correspondingly. It can be seen that the increase in the intensity for the lowest electron energy provides an improvement in the number of created pairs by two orders of magnitude. A pronounced rate growth is also visible, however to a smaller extent, for $E_{0}=5$ GeV and $E_{0}=10$ GeV. The steepest increase occurs up to $\xi\approx 150$, from where on the rate growth becomes significantly flatter.
@@
The S-model would lead to a strong overestimation of the rate practically throughout the whole range of considered laser intensities. This is illustrated by the red dashed line for $E_{0}=2.5$ GeV. It approximately agrees with the D-model rate up to $\xi\approx 70$, which is associated with a maximal $\kappa\approx 2.1$, and strongly deviates for the higher intensities. The deviations to the D-model prediction would be even larger for the higher incident electron energies, which are therefore not included in the figure.
@@
At this point we emphasize that the $\kappa$ values, for which the S- and D-models coincide, are lower for the spectrally resolved pair yields (see discussion of Figs. 3 and 6) than for the total pair yields. As already mentioned before this is because one integrates over the spectrally resolved curve in order to get the total pair yield, thus including also those parts of the bremsstrahlung spectrum where both models are in good agreement.
@@
The total pair yield in an envisaged experiment can be obtained by taking the number of interacting $\gamma$-photons into account. As mentioned above, we assume that the incident electrons come in bunches of 10 pC total charge and that $1\%$ of them will emit bremsstrahlung. Thus in the range $\xi\in[50,300]$, for $E_{0}=2.5$ GeV, from 0.005 to 0.52 pairs per shot can be expected, depending on the applied intensity. Next, by increasing the incident electron energy to 5 GeV [47], the pair yield can be amplified significantly to 0.03 to 1.02 pairs per shot. Lastly, the amount of 0.11 and up to 1.61 created positrons is achievable in the most optimistic scenario with $E_{0}=10$ GeV. Furthermore if we compare the yield between 2.5 GeV and 5 GeV, we observe a fourfold raise in the number of created particles at $\xi=100$ from 0.055 pairs to 0.19 pairs per shot. At $\xi=250$, doubling the energy from 2.5 GeV to 5 GeV still results in approximately doubling the yield, increasing from 0.40 pairs per shot to 0.83 pairs per shot. For 10 GeV we observe 0.44 pairs at $\xi=100$ and up to 1.37 created pairs for $\xi=250$. Hence, the numbers presented here drastically outperform the predictions for ’discovery experiments’ planned in the nearest future (see Sec. III.1) and would thus allow for precision tests of the underlying theory.
@@
At this point, we can quantify the appropriateness of the approximation in Eq. (11) where the longitudinal offset in the $\gamma$-photon trajectory was set to zero, $\tilde{z}=0$. It means that the considered $\gamma$-photons reach the center of the laser focus at time $t=0$, where the laser field strength is maximum. When the trajectory is taken more generally as $\vec{r}(t)=-t\hat{e}_{z}+\tilde{z}+b\hat{e}_{\rho}$, those $\gamma$-photons with $\tilde{z}>0$ ($\tilde{z}<0$) hit the laser pulse maximum by $\tilde{t}=\tilde{z}/2$ earlier (later). An average over the relevant range of $\tilde{z}$-values would need to be taken in a more accurate treatment to calculate the pair yield from a bremsstrahlung $\gamma$-burst. The dependence of the total number of produced pairs per laser shot on the longitudinal offset is depicted in Fig. 9 for $0\leq\tilde{z}\leq z_{R}$. One can see that the curves are very flat, implying that it is not strictly necessary to take an average; accordingly, the simplifying assumption $\tilde{z}=0$ in Eq. (11) works well in the considered range of parameters. The largest relative differences arise for $E_{0}=2.5$ GeV and $\xi=100$, where the pair yield is reduced by about 2.5% (10%) when $\tilde{z}=z_{R}/2$ ($\tilde{z}=z_{R}$); for $E_{0}=10$ GeV and $\xi=250$ the pair yield is instead increased by about 3% (12%) for the same values of $\tilde{z}$. The approximation works well for both large values of $\xi$ as well as large values of $E_{0}$.
@@
is kept constant and the parameter $x$ is varied between 0.5 and 6.5, leading to higher intensity accompanied by narrower beam waist and visa versa. Accordingly, the scaled intensity parameter $\xi=250/x$ in the range $[39,500]$ with the corresponding beam radius $w_{0}=2x\ \mu$m in the intevall $[1\ \mu\text{m},13\ \mu\text{m}]$ was studied for the incident electron energies $E_{0}=2.5$ GeV (red), $E_{0}=5$ GeV (blue) and $E_{0}=10$ GeV (black). Additionally the prediction by the S-Model is displayed for $E_{0}=2.5$ GeV (red dashed) for comparison. Starting from $x\approx 3.5$, corresponding to a maximal $\kappa\approx 2.1$, both models closely coincide with each other.
@@
Figure 10: Number of pairs as a function of the degree of laser focusing obtained by setting $\xi=250/x$ and $w_{0}=2x\ \mu$m and keeping the laser energy constant. The black, blue and red solid lines represent different values of the incident electron energy $E_{0}\in\{2.5,5,10\}$ GeV using the D-model, while the dashed line results from the S-model for $E_{0}=2.5$ GeV.

### v2

@@ III.1 Comparison between S-model and D-model at moderately high field strengths
Figure 3: Comparison between the D-model (solid) and S-model (dashed) results for the differential number of created pairs from different bremsstrahlungs photon energies at $\xi=70$ and $E_{0}=2.5$ GeV.
Figure 3 shows the differential number of created pairs (per radiating electron) in dependence on the normalized bremsstrahlung-photon energy. The set of parameters of a planned experiment on strong-field Breit-Wheeler pair production [13, 33] is considered here: $E_{0}=2.5$ GeV, $\xi=70$ (corresponding to an intensity of $I\approx 10^{22}$ W/cm2), and $\tau=30$ fs. The highest value of the quantum nonlinearity parameter resulting in this scenario is $\kappa\approx 2$ for $f\approx 1$. As the figure displays, the contribution to the pair production yield grows monotoneously with increasing $f$. Up to $f\approx 0.5$ (corresponding to $\kappa\approx 1$), the predictions of the S- and D-models closely agree with each other, but start to slightly deviate afterwards. At the bremsstrahlung endpoint $f\approx 1$, the S-model overestimates the pair yields by about 20%. [^2] We recall that, at $f=1$, the differential number $dN/df$ would actually sharply drop to zero. However, this feature of a real bremsstrahlung spectrum is not covered by the analytical thin-target formula (5), as was explained in Sec. II.1.
@@
Apart from the value of the quantum nonlinearity parameter $\kappa$, we emphasize that also the duration of the interaction influences the relation between S- and D-model predictions. Let us demonstrate this by a simple estimate. In the tunneling-like regime of strong-field Breit-Wheeler pair production at $\kappa\ll 1$, the process rate is given by an analytical expression of the form $R\sim\frac{\alpha m^{2}\kappa}{\omega^{\prime}}\,{\rm e}^{-8/(3\kappa)}$. This compact formula still holds approximately for $\kappa\sim 1$ (see, e.g., Fig.5 in [33]). Taking $\kappa=2$ and $\omega^{\prime}=2.5$ GeV, the rate becomes $R\sim 10^{-7}m$. As a consequence, the time-integrated rate remains smaller than unity for laser pulse durations $\tau\lesssim 10$ fs. Therefore, the S-model is applicable for few-cycle optical laser pulses with duration of the order or below $\approx 10$ fs in this case.
@@ III.2 Pair production in the overcritical field regime
In this subsection, we utilize our D-model and apply Eq. (15) to obtain several important insights into the Breit-Wheeler pair creation process in the overcritical regime with $1<\kappa\lesssim 30$. Also the contributions from the second generation of created pairs will be considered. In contrast to the previous Sec. III.1, here the attention is put on high laser intensities above $10^{22}\ \mathrm{W/cm^{2}}$. These values are achievable in the foreseeable future at the laser facilities such as ELI-NP in Romania, where the investigation of the Breit-Wheeler pair production is firmly embedded into the planning [15]. In the following, we shall therefore use a laser pulse duration of $\tau=20$ fs, as a typical value envisaged at this facility.
@@ III.2.1 Spectrally resolved pair yields
Figure 5: Differential number of created pairs from different bremsstrahlung photon energies at incident electron energy $E_{0}=2.5$ GeV for different values (from bottom to top) of the laser intensity parameter $\xi\in\{100,150,200,300\}$ as indicated. The dashed lines show the D-model results for the first generation of pairs; the solid lines take the additional contributions from the second pair generation into account.
By utilizing the D-model we can analyse the contribution to the number of created pairs from different $\gamma$-photon energies within the bremsstrahlung spectrum. These contributions are depicted in Fig. 5 for different values of $\xi$ between 100 and 300 at a fixed incident electron energy of $E_{0}=2.5$ GeV. The dashed lines refer to the first pair generation, whereas the solid lines show the sum of the first and second pair generations.
@@
We first discuss the properties of the curves for the first pair generation. As we go to high values of $\xi$ we see a clear contrast in the distributions compared to the case in Fig. 3. For $\xi=100$, we observe as before a strictly monotonic increase in the number of created pairs as $f$ grows. In contrast, for higher values of $\xi$, a maximum emerges, indicating a peak contribution at specific $\gamma$-photon energies. For $\xi=150$, the curve already shows a slight bend, for $\xi=200$, we observe a flat global maximum at $f\approx 0.48$, which develops into a pronounced maximum at $f\approx 0.28$ for $\xi=300$. The formation of those maxima results from three effects which influence the creation of first-generation pairs. Firstly, for the parameters under consideration, the local rate of pair creation still approaches to an exponential dependence and thus increases steeply with $\kappa$, which is proportional to the product of $\xi$ and $f$. Thus, for $\xi\lesssim 100$, the exponential rate damping is still strong and therefore the number of pairs grows as $f$ increases, whereas for high values of $\xi$, the field dependence of the rate becomes much flatter (see also Fig. 8 below). Secondly, the number of produced bremsstrahlung photons is much higher for low values of $f$, as depicted in Fig. 2, which starts playing a crucial role in the pair production when the rate growth has left the very steep exponential region and entered into the flatter portion. Lastly, the exponential decay of the bremsstrahlung photons in the D-model is much more pronounced for high values of $\kappa$, which also favors low values of $f$ for the creation of pairs. It is therefore, that we observe the emergence of a maximum at lower values of $f$, because the increase of the rate with higher values of $f$ is counteracted and eventually overcompensated by the falling distribution of bremsstrahlung photons and their decay for high values of $\xi$.
Next we consider the additional contribution from the second generation of pairs. As the solid curves in Fig. 5 show, these contributions increase when $f$ grows. While for $\xi=100$ the enhancement of the pair yield due to the second generation is still marginal, for $\xi=150$, 200 and 300 the enhancement factors at $f\approx 1$ amount to 1.08, 1.19 and 1.49, respectively. As a result of the enhanced pair yields, the solid curve for $\xi=200$ does not feature a maximum but continues to grow slowly. Instead, a pronounced maximum at $f\approx 0.32$ remains in the solid curve for $\xi=300$, which is followed by a shallow minimum at $f\approx 0.68$ and a subsequent distinct increase. Furthermore, from the results for $\xi=200$ and 300 we can infer that the plain D-model of Eq. (14) provides a good approximation to the differential pair yield $dN/df$ up to quantum nonlinearity parameters of $\kappa\approx 6$, where the additional contributions from the second pair generation remain below $\approx 20$%.
@@
A similar observation can be made when we compare the impact of the energy of bremsstrahlung photons for different values of the incident electron energy. In Fig. 6 the contribution to the number of created pairs is plotted for $E_{0}=2.5$ GeV (red), $E_{0}=5$ GeV (blue) and $E_{0}=10$ GeV (black) at fixed $\xi=250$. Maxima in the number of created first-generation pairs (dashed curves) arise at $f\approx 0.35$ for $E_{0}=2.5$ GeV, $f\approx 0.185$ for $E_{0}=5$ GeV, and most distinctly for $E_{0}=10$ GeV at $f\approx 0.0975$. When the second pair generation is taken into account (solid curves), these maxima remain but become less pronounced. Here, the enhancement factors at $f\approx 1$ are 1.33, 1.85 and 2.39 for $E_{0}=2.5$ GeV, 5 GeV and 10 GeV, respectively. As before, the relative contributions from the second pair generation lie below $\approx 20$% up to $\kappa\approx 6$.
@@
For comparison, the S-model prediction for the first pair generation at $E_{0}=2.5$ GeV is represented in Fig. 6 by the dotted red line. It agrees with the D-model result up to $f\approx 0.15$, corresponding to $\kappa\approx 1.1$, but sizably overestimates the yield for higher bremsstrahlung photon energies. In particular, no maximum emerges along the S-model curve. The S-model results for the higher electron energies are not shown in Fig. 6 because they would deviate even more strongly from the D-model. We note that these curves would feature maxima at larger $f$ values because of the counteracting trends between the increase of the rate and the decrease of the bremsstrahlung spectrum. An additional effect contained in the D-model description is the attenuation of the $\gamma$-beam. In result, the maxima of the blue and black dashed lines are located at smaller $f$ values than one would expect based on an S-model consideration.
Figure 6: Contributions to the number of created pairs from different bremsstrahlung photon energies at $\xi=250$. The red, blue and black curves (in this order) represent different values of the incident electron energy $E_{0}\in\{2.5,5,10\}$ GeV. The dashed lines show the results for the first generation of pairs obtained from the D-model; the solid lines also include the second generation of pairs. For comparison, the dotted line represents the result for the first pair generation at $E_{0}=2.5$ GeV predicted by the S-model.
@@
As an illustration, Fig. 7 depicts the values of the spacetime-dependent quantum nonlinearity parameter $\kappa(t,\rho,z,f)$ at $\rho=0$ for 10 GeV incident electrons and the integration region given in Eq. (8). In the upper panel the local values of $\kappa$ for very small photon energy with $f=0.05$ is exposed. Here, the number of created particles is expected to be low as $\kappa$ is of the order of one. On the other hand, for $f=0.1$ the values presented in the middle panel overpass the critical value of $\kappa\approx 1$ and the expected pair yield per $\gamma$-photon is significantly increased, as the heavy exponential damping characteristic for the tunneling-like regime at $\kappa\ll 1$ is largely softened here, so that the course of the pair production rate is distinctly flattened. This observation combined with the high number of produced bremsstrahlung photons (see the graph in Fig. 2) is responsible for the emergence of the maximum in Fig. 6. Lastly, the quantum nonlinearity parameter for $f=0.2$ is shown in the lower panel of Fig. 7. Here, $\kappa\approx 6$ is reached leading to a high rate of the particle creation. However, as the number of bremsstrahlung $\gamma$’s is much lower than for $f=0.1$ (see Fig. 2) and their attenuation in the process is stronger, the resulting outcome for particle production is reduced.
@@ III.2.2 Total pair yields
Figure 8: Dependence of the number of created pairs on the laser intensity for various incident electron energies, color-coded with red, blue and black lines for $E_{0}=2.5$, 5, and 10 GeV, respectively. The dashed lines show the contribution from the first generation of pairs using the D-model; the solid lines additionally include the second generation of pairs. For comparison, the dotted line represents the S-model prediction for the first pair generation at $E_{0}=2.5$ GeV.
@@
In Fig. 8 the total number of pairs produced by a single radiating electron is considered in the intensity range $\xi\in[50,300]$. Moreover, three different values of the incident electron energy $E_{0}\in\{2.5,5,10\}$ GeV are studied and depicted as red, blue and black dashed (solid) lines for the first (first + second) pair generation, correspondingly. It can be seen that the increase in the intensity for the lowest electron energy provides an improvement in the number of created pairs by two orders of magnitude. A pronounced rate growth is also visible, however to a smaller extent, for $E_{0}=5$ GeV and $E_{0}=10$ GeV. The steepest increase occurs up to $\xi\approx 150$, from where on the rate growth becomes significantly flatter.
@@
The S-model prediction for the first pair generation would lead to a strong overestimation of the rate practically throughout the whole range of considered laser intensities. This is illustrated by the red dotted line for $E_{0}=2.5$ GeV. It approximately agrees with the D-model rate up to $\xi\approx 70$, which is associated with a maximal $\kappa\approx 2.1$, and strongly deviates for the higher intensities. The deviations to the D-model prediction would be even larger for the higher incident electron energies, which are therefore not included in the figure.
@@
At this point we emphasize that the $\kappa$ values, for which the S- and D-models coincide, are lower for the spectrally resolved pair yields (see discussion of Figs. 3 and 6) than for the total pair yields. As already mentioned before this is because one integrates over the spectrally resolved curve in order to get the total pair yield, and thus includes also those parts of the bremsstrahlung spectrum where both models are in good agreement.
@@
The total pair yield (including the second generation) in an envisaged experiment can be obtained by taking the number of interacting $\gamma$-photons into account. As mentioned above, we assume that the incident electrons come in bunches of 10 pC total charge and that $1\%$ of them will emit bremsstrahlung. Thus in the range $\xi\in[50,300]$, for $E_{0}=2.5$ GeV, from 0.005 to 0.61 pairs per shot can be expected, depending on the applied intensity. Next, by increasing the incident electron energy to 5 GeV [53], the respective pair yields can be amplified significantly to 0.03 and to 1.44 pairs per shot. Lastly, the amount of 0.12 (at $\xi=50$) and up to 2.81 (at $\xi=300$) created pairs is achievable in the most optimistic scenario with $E_{0}=10$ GeV. For comparison, the yields amount to 0.52, 1.02 and 1.61 first-generation pairs per shot at $\xi=300$ and $E_{0}=2.5$ GeV, 5 GeV and 10 GeV, respectively.
Given the enhancement due to the second pair generation for large values of $E_{0}$ and $\xi$, the question arises whether also contributions from the third pair generation are relevant. It has been shown in [51] that inclusion of the second generation is sufficient for quantum nonlinearity parameters up to $\kappa\approx 16$–20. We have checked that, in our case, the contribution to the first generation from values of $\kappa\geq 16$ is about 8% for the largest parameter combination of $E_{0}=10$ GeV and $\xi=300$. Hence we may conclude that, in the parameter range under consideration, the third pair generation would give only minor additional contributions to the total pair yields.
Continuing our analysis of Fig. 8 by comparing the yield between 2.5 GeV and 5 GeV, we observe an almost fourfold raise in the number of created (first + second generation) particles at $\xi=100$ from 0.056 pairs to 0.20 pairs per shot. At $\xi=250$, doubling the energy from 2.5 to 5 GeV still results in increasing the yield from 0.44 pairs to 1.09 pairs per shot. For 10 GeV we observe 0.49 pairs at $\xi=100$ and up to 2.18 created pairs for $\xi=250$. Hence, the numbers presented here drastically outperform the predictions for ’discovery experiments’ planned in the nearest future (see Sec. III.1) and would thus allow for precision tests of the underlying theory.
The plain D-model of Eq. (15) provides a good approximation to the total pair yields in Fig. 8 up to maximal quantum nonlinearity parameters of $\kappa\approx 12$–15, where the relative additional contributions from the second pair generation remain below $\approx 20$%. These $\kappa$-values are distinctly larger than those found in the discussion of the spectrally resolved pair yields $dN/df$ in Figs. 5 and 6. Similarly as before when comparing the S- and D-model predictions, this is because the total pair yield involves an integration over the whole bremsstrahlung spectrum, which contains $f$-regions where the second pair generation does not give sizeable contributions.
At this point, we can quantify the appropriateness of the approximation in Eq. (11) where the longitudinal offset in the $\gamma$-photon trajectory was set to zero, $\tilde{z}=0$. It means that the considered $\gamma$-photons reach the center of the laser focus at time $t=0$, where the laser field strength is maximum. When the trajectory is taken more generally as $\vec{r}(t)=-t\hat{e}_{z}+\tilde{z}+b\hat{e}_{\rho}$, those $\gamma$-photons with $\tilde{z}>0$ ($\tilde{z}<0$) hit the laser pulse maximum by $\tilde{t}=\tilde{z}/2$ earlier (later). An average over the relevant range of $\tilde{z}$-values would need to be taken in a more accurate treatment to calculate the pair yield from a bremsstrahlung $\gamma$-burst. The dependence of the total number of pairs per laser shot on the longitudinal offset is depicted in Fig. 9 for $0\leq\tilde{z}\leq z_{R}$. Here, the first generation of produced pairs is considered. One can see that the curves are very flat, implying that it is not strictly necessary to take an average; accordingly, the simplifying assumption $\tilde{z}=0$ in Eq. (11) works well in the considered range of parameters. The largest relative differences arise for $E_{0}=2.5$ GeV and $\xi=100$, where the pair yield is reduced by about 2.5% (10%) when $\tilde{z}=z_{R}/2$ ($\tilde{z}=z_{R}$); for $E_{0}=10$ GeV and $\xi=250$ the pair yield is instead increased by about 3% (12%) for the same values of $\tilde{z}$. The approximation works well for both large values of $\xi$ as well as large values of $E_{0}$.
@@
is kept constant and the scaling parameter $x$ is varied between 1 and 6.5, leading to higher intensity accompanied by narrower beam waist and visa versa. Accordingly, the scaled intensity parameter $\xi=250/x$ in the range $[39,500]$ with the corresponding beam radius $w_{0}=2x\ \mu$m in the intevall $[2\,\mu\text{m},13\,\mu\text{m}]$ was studied for the incident electron energies $E_{0}=2.5$ GeV (red), $E_{0}=5$ GeV (blue) and $E_{0}=10$ GeV (black).
@@
Figure 10: Number of pairs as a function of the degree of laser focusing obtained by setting $\xi=250/x$ and $w_{0}=2x\ \mu$m and keeping the laser energy constant. The red, blue and black colors refer to different values of the incident electron energy $E_{0}\in\{2.5,5,10\}$ GeV. The dashed lines show the results for the first generation of pairs using the D-model; the solid lines include in addition the second generation of pairs.